For the reaction N20g) 3H2(g) 2 NH3(g) AGO 23.6 kJ and AS 198.7 J/K at 345 K and 1 atm This reaction is (reactant, product) favored under standard conditions at 345 K. The standard enthalpy change for the reaction of 2.30 moles of N20g) at this temperature would be For the reaction 2 H2020) 2 H20(l) o2 (g) Go 236.9 kJ and AHO 196.00 kJ at 325 K and 1 atm This reaction is (reactant, product) favored under standard conditions at 325 K. The entropy change for the reaction of 1.57 moles of H2020 at this temperature would be For the reaction Fe203 (s) 3H20g 2Fe(s) 3H20(g) Ho 98.8 kJ and ASO 142.5 J/K The standard free energy change for the reaction of 2.35 moles of Fe203(s) at 281 K, 1 atm would be This reaction is (reactant, product) favored under standard conditions at 281 K. Assume that AHo and AS are independent of temperature. kJ. J/K.
Answer #1

Q1.

since dG is negative, this is favoured towad the right

so this reaction is product favoued

for

dG = dH - T*dS

dH = dG + T*dS = -23.6 + 345*-(0.1987)

dH = -92.15 kJ/mol

for 1 mol of N2 --> -92.15

for 2.3 mol --> -92.15*2.3 = -211.945 kJ will be formed

Q2

This is once again -dG, so this is favoured toward products

Find Entopy given:

dG = dH - T*dS

dS = (dG-dH)/(-T)

dS = (-236.9 - -196)/(-325) = 0.1258 kJ/molK

this is per 2 mol of H2O2

for

1.57 moles --> 1.57/2*0.1258 = 0.098753 kJ/K = 0.098753*10^3 J/K = 98.75 J/K

Q3.

finally

dG = dH - T*dS = 98.8 -281*(0.1425) = 58.757 kJ

this must favour reactants, since it is positive with respect to dG value

So

dG when 1 mol reat --> 58.757 kJ

dG when 2.35 moles react --> 58.757*2.35 = 138.07 kJ